[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Apr 04, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Apr 4 16:27:52 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
Please continue to report NAWIPS problems in the TAFB shift log so they can  
be included in the tracking spreadsheet mentioned below.

Monica added Chris Burr to the TAFB shift log. She also removed Trisha and  
Jessica while she was working on that.

Forecasters were unable to retrieve AWIPS and GFE windows on the LX2  
workstation after the windows are minimized. GFE closed completely after it  
was minimized. Eric C. reported the same problem on the mid shift. Chris  
Mello will look into that.

The NCO visitors will be working with Monica and Chris Mello on AWIPS use  
cases the rest of the week. They will need to shadow TAFB forecasters on  
some shifts as part of that work.
HSU:

Other items:
Matt Sardi prepared a Goolge doc with remaining issues related to the RHEL6  
conversion of the NAWIPS systems. Paul Obe from NCO is busy with the NetApp  
installation the rest of today, but we will coordinate with him on a plan  
to address the items on the spreadsheet over the next several days.

The 12Z GFS was more than 2 hours late yesterday, but that was probably a  
result of some network disruption due to the installation of the new  
NetApp. However, since the network change by NCO 2 weeks ago, the GFS has  
routinely been taking 20 to 40 min longer to complete. We were getting  
bandwidth we were not supposed to get, but NCO fixed that "error". NCO is  
aware of the slowness issue now that our network is configured properly,  
and is working to double our bandwidth before the start of the hurricane  
season.


NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170404/fe3ebdf8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list